OpenStudy (anonymous):

y - y1 = m(x - x1) HELP (64,65) (72,71)

OpenStudy (anonymous):

you choose one of those 2 points to plug in the equation. of course choose the one with the smaller numbers which would be (64,65) and then plug it in. y-65=m(x-64) then when you have that done you will need to find the slope which is m= \[\frac{ 71-65 }{ 72-64 }\] what would the slope be?

OpenStudy (anonymous):

were you asked to leave the answer as standard, slope intercept or point slope form?

OpenStudy (anonymous):

i was asked to put it in point slope intercept form

OpenStudy (anonymous):

ok then all you need to do is find the slope (m) and plug it in and that would be your answer. hope it helped!

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

best response?